OpenStudy (anonymous):

x^2-4^2=9 y-2x=0 solve by substitution method..can someone show me the steps

OpenStudy (zepp):

\(y-2x=0\) Isolate y; \(y=2x\)

OpenStudy (zepp):

Is it x^2-4\(\huge y\)^2=9?

OpenStudy (zepp):

If so, substitute y into the first equation you got \(\begin{align} \\ x^2-4y^2&=9 \\ x^2-4(2x)^2&=9 \\ x^2-4(4x^2)&=9 \\ x^2-16x^2 &=9 \\ x^2-16x^2 &=9 \\ -15x^2&=9 \\ x^2&=0.6 \\ x&=\pm \sqrt{0.6} \end{align}\)

OpenStudy (anonymous):

the book says that the answer is none tho

OpenStudy (anonymous):

but i don't how they got that answer

OpenStudy (zepp):

Are you sure the question is \(x^2-4^2=9 \\ y-2x=0\) ?

jimthompson5910 (jim_thompson5910):

zepp you go from -15x^2=9 to x^2=0.6 but you lost the sign somehwere as it should be x^2=-0.6

OpenStudy (anonymous):

yes I am sure

jimthompson5910 (jim_thompson5910):

somewhere*

jimthompson5910 (jim_thompson5910):

so that explains why there are no real answers

OpenStudy (zepp):

Ohh, my minus :( Where are you?? \(\pm \sqrt{-0.6}\) then, therefore no real solution.

OpenStudy (zepp):

but asker states that's \(x^2−4^2=9\) and not \(x^2-4y^2=9\)

jimthompson5910 (jim_thompson5910):

Well if the first equation is x^2−4^2=9, then there are 2 ordered pair real solutions. But the book states that there are no real solutions. So my guess is that it should be x^2−4y^2=9

OpenStudy (anonymous):

the problem is x^2-4^2=9 y-2x=0 i looked again

jimthompson5910 (jim_thompson5910):

then there's a typo in the problem or there's a mistake in the answer given

OpenStudy (zepp):

\(x^2-4^2=9\\x^2-16=9\\x^2 =9+16\\x^2=25\\x=\pm\sqrt{25}\) \(x=5;-5\) \(y-2x=0\\y-2(5)=0\\y-10=0\\y=10\) Ordered pair 1: \((5,10)\) \(y-2x=0\\y-2(-5)=0\\y+10=0\\y=-10\) Ordered pair 1: \((-5,-10)\)

OpenStudy (zepp):

There must be a typo in the book :)

OpenStudy (anonymous):

ok so, the last thing u sent me is how u solve the problem by the substitution method?

OpenStudy (zepp):

Yes, I found x, and substituted x into the second equation to find y.

OpenStudy (anonymous):

ok it, thanks!

OpenStudy (zepp):

You are welcome :)